A triangle has a perimeter of 30m. The length of the 2nd side is 5m more than the length of the 1st side. The third side is 3 meters less than twice the 1st side. How long is each side?
These always give me trouble!
My solution came out with two negative answers so I know they cannot be right. How would I set up the equation?
Here is what I have:
P=30m
2nd side=P+5m
3rd side=30m(2)-3m
30m=P+5m+30m(2)-3m
30m=P+5m+60m-3m
30m=P+62m
30m-62m=P+62m-62m
-32m=P
HELP PLEASE!!




Answer by stanbon(75887)

A triangle has a perimeter of 30m.
The length of the 2nd side is 5m more than the length of the 1st side.
The third side is 3 meters less than twice the 1st side.
How long is each side?
---
Let 1st side be "x"
Then 2nd side is "x+5"
And 3rd side is "2x-3"
----------------------------
Equation:
x + x+5 + 2x-3 = 30
4x + 2 = 30
4x = 28
------------------
x = 7m (1st side)
x+5 = 12m (2nd side)
2x-3 = 11m (3rd side)
==========================
Cheers,
Stan H.
=============